OpenStudy (anonymous):

Eliminate the parameter. x = t - 3, y=2/t+5

OpenStudy (anonymous):

I honestly have no idea how to even start this

OpenStudy (anonymous):

All your doing is solving for t and plugging it into the other equation so you get rid of the t, just a simple elimination problem.

OpenStudy (anonymous):

so like x = t - 3 t = x + 3 y = 2 / x + 8

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Yeah